OpenStudy (anonymous):

To which graph does the point (2, -3) belong? Help Please What is the equation in standard form of the line which passes through (-2, -3) and has a slope of ?

OpenStudy (anonymous):

http://openstudy.com/users/jamesj

OpenStudy (anonymous):

slope ???

OpenStudy (anonymous):

the slope for the second question is 1/2

OpenStudy (anonymous):

use this formula y-y1 = m(x-x1)

OpenStudy (anonymous):

can you just show me the steps

OpenStudy (anonymous):

ok put y1 = -3 , x1 = -2 and m = 1/2 in the formula i gave you

OpenStudy (anonymous):

ya but what is the answer

OpenStudy (anonymous):

what you got?? first show me that then i will

OpenStudy (anonymous):

alright one sec

OpenStudy (anonymous):

y-(-2)=1/2(x-(-2)

OpenStudy (anonymous):

y-(-3) = 1/2(x-(-2)) now go one step more

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

y-(-3)=1/2x-(-1)

OpenStudy (anonymous):

samtasticc can you help with the first one